### v1

@@ 6      Initialize $\hat{z}^{t}_{1}=\mathcal{Q}(x_{r_{t}:t})$
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If there exists a finite sequence of uniformly noisily generatable classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$, then $\mathcal{H}$ is noisily generatable in the limit.
@@
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Suppose there exists a finite sequence of uniformly noisily generatable classes $\mathcal{H}_{1},\mathcal{H}_{2},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$. By Theorem 3.1, we know that for all $i\in[k]$, we have that
$\Bigl{|}\bigcap_{h\in\mathcal{H}_{i}}\operatorname{supp}(h)\Bigl{|}=\infty.$
For every $i\in[k]$, let $z^{i}_{1},z^{i}_{2},\dots$ denote the elements of $\bigcap_{h\in\mathcal{H}_{i}}\operatorname{supp}(h)$ sorted in their natural ordering. Consider the following generator $\mathcal{G}$. Given any noisy enumeration $x_{1},x_{2},\dots$ and any $t\in\mathbb{N}$, $\mathcal{G}$ first computes
for every $i\in[k]$. Then, $\mathcal{G}$ computes $i_{t}=\operatorname*{arg\,max}_{i\in[k]}p^{i}_{t}.$ Finally, $\mathcal{G}$ plays arbitrarily from $\{z^{i_{t}}_{1},z^{i_{t}}_{2},\dots\}\setminus\{x_{1},\dots x_{t}\}.$ We claim that $\mathcal{G}$ generates from $\mathcal{H}$ in the limit.
To see why, let $h\in\mathcal{H}$ and $x_{1},x_{2},\dots$ be the hypothesis and noisy enumeration selected by the adversary. Let $i^{\star}\in[k]$ be such that $h\in\mathcal{H}_{i^{\star}}$ and $s^{\star}\in\mathbb{N}$ be such that for all $t\geq s^{\star}$, we have that $x_{t}\in\operatorname{supp}(h)$. Such an $s^{\star}$ exists because $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty.$ Let $S^{\star}\subseteq[k]$ be such that $i\in S^{\star}$ if and only if $\{z^{i}_{1},z^{i}_{2},\dots\}\subseteq\{x_{1},x_{2},\dots\}.$ Note that $i^{\star}\in S^{\star}$ by definition of $z^{i^{\star}}_{1},z^{i^{\star}}_{2},\dots$ and the fact that $x_{1},x_{2},\dots$ is a noisy enumeration. By definition of $S^{\star}$ and $p_{t}^{i}$ it must be the case that for all $j\notin S^{\star}$, we have that $\sup_{t\in\mathbb{N}}p_{t}^{j}<\infty.$ Since $k<\infty$, we then get that $\max_{j\notin S^{\star}}\sup_{t\in\mathbb{N}}p_{t}^{j}<\infty.$ On the other hand, for all $i\in S^{\star}$, we have that $p_{t}^{i}\rightarrow\infty.$ More precisely, for every $i\in S^{\star}$, there exists a finite $t_{i}\in\mathbb{N}$ such that $p^{i}_{t_{i}}\geq\max_{j\notin S^{\star}}\sup_{t\in\mathbb{N}}p_{t}^{j}.$ Since $|S^{\star}|<\infty$, we also have that $t^{\star}:=\max_{i\in S^{\star}}t_{i}<\infty.$ Our proof is complete after noting that for all $t\geq\max\{s^{\star},t^{\star}\}$, we have that $i_{t}\in S^{\star}$ and $\{z^{i_{t}}_{1},z^{i_{t}}_{2},\dots\}\setminus\{x_{1},\dots x_{t}\}\subseteq\operatorname{supp}(h).$
4 Discussion and Open Questions
In this paper, we introduced several notions of noisy generatability. We highlight some important directions of future work that we left open.
Characterizations of Noisy Non-uniform Generatability and Noisy Generatability in the Limit. An important direction of future work is to provide a complete and concise characterization of which classes are noisily non-uniformly generatable and which classes are noisily generatable in the limit.
Question 4.1.
What characterizes noisy non-uniform generatability and noisy generatability in the limit?
Note that Li et al. (2024) also pose the complete characterization of (noiseless) generatability in the limit as an open question, but do provide a complete characterization of (noiseless) non-uniform generatability. We conjecture that the correct characterization of noisy non-uniform generatability will need to go beyond our sufficiency and necessity conditions in Section 3.3.
Noisy Generatability in the Limit via Membership Oracles. In addition to showing that all countable classes are generatable in the limit, Kleinberg and Mullainathan (2024) provide a computable algorithm for doing so when given access to a membership oracle. In our notion, a membership oracle $\mathcal{O}_{\mathcal{H}}:\mathcal{H}\times\mathcal{X}\rightarrow\{0,1\}$, takes as input a hypothesis $h\in\mathcal{H}$ and an example $x\in\mathcal{X}$ and outputs $\mathbbm{1}\{h(x)=1\}$. Although we show that all countable classes are also noisily generatable in the limit, the focus of this paper was mainly information-theoretic in nature. This motivates our next open question.
Question 4.2.
Given access to a membership oracle, does there exists a computable algorithm that can noisily generate in the limit from any countable class?
Noiseless vs. Noisy Generatability In Section 3.2, we showed that noiseless uniform generatability can be strictly easier than noisy uniform generatability. Our third open question asks whether this is also true for non-uniform generatability.
Question 4.3.
Is there a class which is (noiseless) non-uniformly generatable but not noisily non-uniformly generatable?
Recall that in Section 3.3, we showed that all countable classes are noisily non-uniformly generatable. Thus, an affirmative resolution to Question 4.3 must involve an uncountably infinite class. Nevertheless, we conjecture that such a separation is true given the separation between noisy and noiseless uniform generation. One can also ask this question with regards to generatability in the limit.
Question 4.4.
Is there a class which is (noiseless) generatable in the limit but not noisily generatable in the limit?
Again, a positive affirmation to this questions must involve an uncountably infinite class.
Appendix A The Hardness of Non-uniform Noisy Generatability
@@
Definition A.1 (Non-uniform Noisy Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is non-uniformly noisily generatable, if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, there exists $d^{\star}\in\mathbbm{N}$ such that for any sequence $x_{1},x_{2},\dots$ with
@@
Unfortunately, non-uniform noisy generatability is still a restrictive property as not all finite classes are non-uniformly noisily generatable. This is in contrast to noisy uniform generatability as shown by Corollary 3.3.
Lemma A.1 (Not All Finite Classes are Non-uniformly Noisily Generatable).
Let $\mathcal{X}$ be countable. There exists a finite class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ which satisfies the UUS property that is not non-uniformly noisily generatable.
@@
Let $\mathcal{X}=\mathbbm{N}$ and consider the class =$\mathcal{H}=\{h_{e},h_{o}\}$ such that $h_{e}(x)=\mathbbm{1}\{x\text{ is even}\}$ and $h_{o}(x)=\mathbbm{1}\{x\text{ is odd}\}.$ We will show that $\mathcal{H}$ is not non-uniformly noisily generatable. We need to show that for every generator $\mathcal{G}$, there exists a hypothesis $h\in\mathcal{H}$ such that for every $d\in\mathbb{N}$, there exists a sequence $x_{1},x_{2},\dots$ with

### v2

@@ 3.4 Noisy Generatability in the Limit
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If there exists a finite sequence of uniformly noise-independent generatable classes $\mathcal{H}_{1},\dots,\mathcal{H}_{k}$ such that $\mathcal{H}=\bigcup_{i=1}^{k}\mathcal{H}_{i}$, then $\mathcal{H}$ is noisily generatable in the limit.
The proof of Theorem 3.10 is similar to the proof Theorem 3.10 from Li et al. (2024) and so we defer it to Appendix G.
4 Discussion and Open Questions
In this paper, we introduced several notions of noisy generatability and made progress towards characterizing which classes are noisily generatable. We highlight some important directions of future work.
Characterizations of Non-uniform Noise-dependent Generatability and Noisy Generatability in the Limit. An important direction of future work is to provide a complete and concise characterization of which classes are non-uniformly noise-dependent generatable and noisily generatable in the limit. Note that Li et al. (2024) also pose the complete characterization of (noiseless) generatability in the limit as an open question, but do provide a complete characterization of (noiseless) non-uniform generatability. We conjecture that the correct characterization of non-uniform noise-dependent generatability will need to go beyond our sufficiency and necessity conditions in Section 3.3.
Noisy Generatability in the Limit via Membership Oracles. In addition to showing that all countable classes are generatable in the limit, Kleinberg & Mullainathan (2024) provide a computable algorithm for doing so when given access to a membership oracle. A membership oracle, $\mathcal{O}_{\mathcal{H}}:\mathcal{H}\times\mathcal{X}\rightarrow\{0,1\}$, takes as input a hypothesis $h\in\mathcal{H}$ and an example $x\in\mathcal{X}$ and outputs $\mathbbm{1}\{h(x)=1\}$. Although we show that all countable classes are also noisily generatable in the limit, the focus of this paper was information-theoretic in nature. This motivates our next open question. Given access to a membership oracle, does there exists a computable algorithm that can noisily generate in the limit from any countable class?
Acknowledgements
The authors thank anonymous reviewer gxU2 for catching a bug in Theorem 3.1 and providing a fix. VR acknowledges support from the NSF Graduate Research Fellowship Program (GRFP).
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Definitions of Noiseless Generation
Definition A.1 (Uniform Generatability (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly generatable, if there exists a generator $\mathcal{G}$ and $d^{\star}\in\mathbb{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h)$, if there exists $t^{\star}\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Definition A.2 (Non-uniform Generatability (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is non-uniformly generatable if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, there exists a $d^{\star}\in\mathbb{N}$ such that for any sequence $x_{1},x_{2},\dots$ with $\{x_{1},x_{2},\dots\}\subseteq\operatorname{supp}(h)$, if there exists $t^{\star}\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Definition A.3 (Generatability in the Limit (Li et al., 2024)).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is generatable in the limit if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, and any enumeration $x_{1},x_{2},\dots$ of $\operatorname{supp}(h)$, there exists a $t^{\star}\in\mathbb{N}$ such that $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Appendix B Summary of Results for Noiseless Generation
In the noiseless setting, Kleinberg & Mullainathan (2024) initiated the study of generation in the limit by showing that all countable classes are generatable in the limit and that all finite classes are uniformly generatable. Following this result, Li et al. (2024) provided a complete characterization of uniform generation in terms of a new combinatorial parameter termed the Closure dimension.
Definition B.1 (Closure dimension (Li et al., 2024)).
The Closure dimension of $\mathcal{H}$, denoted $\operatorname{C}(\mathcal{H})$, is the largest natural number $d\in\mathbb{N}$ for which there exists distinct $x_{1},\dots,x_{d}\in\mathcal{X}$ such that $\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},0}\neq\bot$ and $|\langle x_{1},\dots,x_{d}\rangle_{\mathcal{H},0}|<\infty.$ If this is true for arbitrarily large $d\in\mathbb{N}$, then we say that $\operatorname{C}(\mathcal{H})=\infty.$ On the other hand, if this is not true for $d=1$, we say that $\operatorname{C}(\mathcal{H})=0.$
Theorem B.2 (Theorem 3.3. in Li et al. (2024)).
Theorem B.3 (Theorem 3.5 in Li et al. (2024)).
A class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ is non-uniformly generatable if and only if there exists a non-decreasing sequence of classes $\mathcal{H}_{1}\subseteq\mathcal{H}_{2}\subseteq\dots$ such that $\mathcal{H}=\bigcup_{i=1}^{\infty}\mathcal{H}_{i}$ and $\operatorname{C}(\mathcal{H}_{i})<\infty$ for every $i\in\mathbb{N}$.
As a corollary, Li et al. (2024) establish that all countable classes are actually non-uniformly generatable. Charikar & Pabbaraju (2024) also established this result independently. With regards to generatability in the limit, Kleinberg & Mullainathan (2024) prove that all countable classes are generatable in the limit. Li et al. (2024) provide an alternate sufficiency condition in terms of the Closure dimension.
Theorem B.4 (Theorem 3.10 in Li et al. (2024)).
Appendix C An Alternate Version of Uniform Noise-independent Generatability
In this section, we consider an alternate (weaker) version of uniform noise-independent generatability by measuring the sample complexity only in terms of the valid positive examples observed.
Definition C.1 (Alternate Uniform Noise-Independent Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noise-independent generatable, if there exists a generator $\mathcal{G}$ and $d^{\star}\in\mathbb{N}$, such that for every $h\in\mathcal{H}$ and any sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty$, if there exists $t^{\star}\in\mathbb{N}$ such that |$|\{x_{1},\dots,x_{t^{\star}}\}\cap\operatorname{supp}(h)|=d^{\star}$, then $\mathcal{G}(x_{1:s})\in\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$ for all $s\geq t^{\star}$.
Note that the only difference between Definition C.1 and D.1 is how $d^{\star}$ is measured. In Definition C.1, $d^{\star}$ captures only the number of distinct positive examples in the stream, while $d^{\star}$ in Definition D.1 measures the total number of distinct examples in the stream. We start off by proving a simple necessity condition for our alternate notion of uniform noise-independent generatability.
Lemma C.2 (Necessary condition for Alternate Uniform Noise-independent Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. If there exists a subclass $\mathcal{F}\subseteq\mathcal{H}$ and a hypothesis $f\in\mathcal{F}$ such that $\left|\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)\right|<\infty$ and $\left|\bigcap_{h\in\mathcal{F}\setminus\{f\}}\operatorname{supp}(h)\right|=\infty$, then $\mathcal{H}$ is not uniformly noise-independent generatable according to Definition C.1.
Like Theorem 3.1, Lemma C.2 is a hardness result – it shows that finite classes with just two hypotheses may not be uniformly noise-independent generatable even when the sample complexity is measured in terms of positive example.
@@ Appendix C An Alternate Version of Uniform Noise-independent Generatability
(of Lemma C.2) Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Let $\mathcal{F}\subseteq\mathcal{H}$ be any subset of $\mathcal{H}$ such that $\Bigl{|}\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)\Bigl{|}<\infty$ and for which there exists an $f\in\mathcal{F}$ such that $\Bigl{|}\bigcap_{h\in\mathcal{F}\setminus\{f\}}\operatorname{supp}(h)\Bigl{|}=\infty.$ We need to show that $\mathcal{H}$ is not uniformly noise-independent generatable. We will show that $\mathcal{F}$ is not uniformly noise-independent generatable which will imply that $\mathcal{H}$ is not uniformly noise-independent generatable. To that end, we need to show that for every $\mathcal{G}$ and any $d\in\mathbb{N}$, there exists a $h\in\mathcal{F}$ and a sequence $x_{1},x_{2},\dots$ with $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h)\}<\infty,$ such that for every $t\in\mathbb{N}$ where |$|\{x_{1},\dots,x_{t}\}\cap\operatorname{supp}(h)|=d$, there exists an $s\geq t$ such that $\mathcal{G}(x_{1:s})\notin\operatorname{supp}(h)\setminus\{x_{1},\dots,x_{s}\}$. To that end, fix a generator $\mathcal{G}$ and a number $d\in\mathbb{N}$. Let $x_{1},\dots,x_{d}$ be a sequence of $d$ unique examples in $\bigcap_{h\in\mathcal{F}\setminus\{f\}}\operatorname{supp}(h)$. Let $z_{1},\dots,z_{d}$ be any $d$ unique examples in $\operatorname{supp}(f)\setminus\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)$ and consider the sequence of $2d$ unique examples obtained by concatenating $z_{1},\dots,z_{d}$ to the end of $x_{1},\dots,x_{d}$. Denote this new sequence by $x_{1:2d}.$ Let $q=|\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)|$ and let $v_{1},\dots,v_{q}$ be its elements sorted in its natural order. Concatenate $v_{1},\dots,v_{q}$ to the end of $x_{1:2d}$ and denote this new sequence of unique examples as $x_{1:2d+q}.$ Observe that for every $h\in\mathcal{F}$, we have that $|\operatorname{supp}(h)\cap\{x_{1},\dots x_{2d+q}\}|\geq d.$ Let $\hat{x}_{2d+q}=\mathcal{G}(x_{1},\dots,x_{2d+q})$ and suppose without loss of generality that $\hat{x}_{2d+q}\notin\{x_{1},\dots,x_{2d+q}\}.$ Let $h^{\star}\in\mathcal{F}$ be a be hypothesis such that $\hat{x}_{2d+q}\notin\operatorname{supp}(h^{\star}).$ Such a hypothesis must exist because $\hat{x}_{2d+q}\notin\{x_{1},\dots,x_{2d+q}\}$ and $\bigcap_{h\in\mathcal{F}}\operatorname{supp}(h)\subseteq\{x_{1},\dots,x_{2d+q}\}.$ Let $x_{2d+q+1},x_{2d+q+2},\dots$ be any completion of the stream such that $\{x_{2d+q+t}\}_{t=1}^{\infty}\subseteq\operatorname{supp}(h^{\star})$ and $\{x_{2d+q+t}\}_{t=1}^{\infty}\cap\{x_{t}\}_{t=1}^{2d+q}=\emptyset.$
We are now ready to complete the proof. Let $h^{\star}$ and $x_{1},x_{2},\dots$ be the hypothesis and sequence chosen above. Then, by definition, observe that $\sum_{t=1}^{\infty}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}=\sum_{t=1}^{2d+q}\mathbbm{1}\{x_{t}\notin\operatorname{supp}(h^{\star})\}\leq d<\infty.$ Moreover, we also know that $|\operatorname{supp}(h^{\star})\cap\{x_{1},\dots,x_{2d+q}\}|\geq d$ and the first $2d$ examples are distinct, implying that there exists exactly one time point $t\leq 2d+q$ such that $|\operatorname{supp}(h^{\star})\cap\{x_{1},\dots x_{2d+q}\}|=d.$ Finally, noting that $2d+q\geq t$ and that $\mathcal{G}(x_{1},\dots,x_{2d+q})\notin\operatorname{supp}(h^{\star})\setminus\{x_{1},\dots,x_{2d+1}\}$ completes the proof that $\mathcal{F}$ is not uniformly noise-independent generatable since $\mathcal{G}$ and $d$ were chosen arbitrarily. ∎
Finally, Theorem C.3 provides a full characterization of uniform noise-independent generatability (Definition D.1 in terms of the noisy closure dimension.
Theorem C.3 (Characterization of Uniform Noise-independent Generatability).
Let $\mathcal{X}$ be countable and $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ satisfy the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is uniformly noise-independent generatable if and only if $\sup_{n}(\operatorname{NC}_{n\in\mathbb{N}}(\mathcal{H})-n)<\infty$.
As the proof follows similar techniques as those in the main text, we only provide the proof sketch here.
Proof.
(sketch of Theorem C.3) For the necessity direction, suppose that $\sup_{n}(\operatorname{NC}_{n}(\mathcal{H})-n)=\infty$. Then for every $d\in\mathbb{N}$, we can find a $t\geq d$ and a sequence $x_{1},\dots,x_{t}$, such that $|\langle x_{1},...,x_{t}\rangle_{\mathcal{H},t-d}|<\infty.$ Hence, by padding $x_{1},\dots,x_{t}$ with any remaining elements in $\langle x_{1},...,x_{t}\rangle_{\mathcal{H},t-d}$, we can force the Generator to make a mistake while ensuring that the hypothesis chosen is consistent with at least $d$ examples in the stream. For the sufficiency direction, if $\sup_{n}(\operatorname{NC}_{n}(\mathcal{H})-n)<\infty$, then there exists a $d\in\mathbb{N}$ such that for every $t\geq d$ and distinct $x_{1},...,x_{t}$, we have that either $\langle x_{1},...,x_{t}\rangle_{\mathcal{H},t-d}=\bot$ or $|\langle x_{1},...,x_{t}\rangle_{\mathcal{H},t-d}|=\infty.$ Thus, the algorithm which for $t\geq d$ plays from $\langle x_{1},...,x_{t}\rangle_{\mathcal{H},t-d}\setminus\{{x_{1},\dots,x_{t}\}}$ if $\langle x_{1},...,x_{t}\rangle_{\mathcal{H},t-d}\neq\bot$ is guaranteed to succeed. ∎
Appendix D The Hardness of Non-uniform Noise-independent Generatability
@@
Definition D.1 (Non-uniform Noise-independent Generatability).
Let $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ be any hypothesis class satisfying the $\operatorname{UUS}$ property. Then, $\mathcal{H}$ is non-uniformly noise-independent generatable, if there exists a generator $\mathcal{G}$ such that for every $h\in\mathcal{H}$, there exists $d^{\star}\in\mathbbm{N}$ such that for any sequence $x_{1},x_{2},\dots$ with
@@
Unfortunately, non-uniform noise-independent generatability is still a restrictive property as not all finite classes are non-uniformly noise-independent generatable. This is in contrast to uniform noise-dependent generatability as shown by Corollary 3.4.
Lemma D.2 (Not All Finite Classes are Non-uniformly Noise-independent Generatable).
Let $\mathcal{X}$ be countable. There exists a finite class $\mathcal{H}\subseteq\{0,1\}^{\mathcal{X}}$ which satisfies the UUS property that is not non-uniformly noise-independent generatable.
@@
Let $\mathcal{X}=\mathbbm{N}$ and consider the class =$\mathcal{H}=\{h_{e},h_{o}\}$ such that $h_{e}(x)=\mathbbm{1}\{x\text{ is even}\}$ and $h_{o}(x)=\mathbbm{1}\{x\text{ is odd}\}.$ We will show that $\mathcal{H}$ is not non-uniformly noise-independent generatable. We need to show that for every generator $\mathcal{G}$, there exists a hypothesis $h\in\mathcal{H}$ such that for every $d\in\mathbb{N}$, there exists a sequence $x_{1},x_{2},\dots$ with